A penny is dropped from the observation deck of the Empire State building (369 m above the ground). With what velocity does it strike the ground?


1
Expert's answer
2022-03-08T16:58:57-0500

The final velosity of a coin

"v=\sqrt{2gh}=\sqrt{2*9.8*369}=85\:\rm m/s"